OpenStudy (anonymous):

If the odds of winning a raffle are 2:229, what is the probability of winning? 2/229 229/2 229/231 2/231

OpenStudy (anonymous):

last one.

jimthompson5910 (jim_thompson5910):

P(winning) = 2/(2+229) P(winning) = 2/231

jimthompson5910 (jim_thompson5910):

In general, if the odds in favor are m:n, then P(winning) = m/(m+n)